OpenStudy (surana):

Given a = 2, B = -2, and k =3, what is the equation of the graph if the parent function is y = /-x?

OpenStudy (surana):

The /- is meant to be the square root symbol.

OpenStudy (surana):

And possible answers stand as A: y = 2 /- 2x+3 B: y = -2 /- 2x +3 C: y = 2 /- negative 2x -3 D: Y = 2 /- negative 2x + 3

OpenStudy (triciaal):



OpenStudy (triciaal):